# Night-Rider: Nocturnal Vision-aided Localization in Streetlight Maps Using Invariant Extended Kalman Filtering

### 01 Feb 2024 v1

@@ Abstract
Vision-aided localization for low-cost mobile robots in diverse environments has attracted widespread attention recently. Although many current systems are applicable in daytime environments, nocturnal visual localization is still an open problem owing to the lack of stable visual information. An insight from most nocturnal scenes is that the static and bright streetlights are reliable visual information for localization. Hence we propose a nocturnal vision-aided localization system in streetlight maps with a novel data association and matching scheme using object detection methods. We leverage the Invariant Extended Kalman Filter (InEKF) to fuse IMU, odometer, and camera measurements for consistent state estimation at night. Furthermore, a tracking recovery module is also designed for tracking failures. Experiments on multiple real nighttime scenes validate that the system can achieve remarkably accurate and robust localization in nocturnal environments.
@@ I Introduction
  • The Invariant Extended Kalman Filter (InEKF) fusing IMU, odometer, and visual measurements is derived for the state estimation and localization.
@@ II System Overview
Fig. 2 shows the framework of the proposed nocturnal localization system. The streetlight map is constructed using only the 3D streetlight points. The system fuses monocular camera, IMU and odometer measurements and leverages the right InEKF [15] for accurate and consistent state estimation according to the right-invariant observation models of above sensors[16, 17, 18]. Specifically, the robot state is propagated using IMU measurements, and the odometer measurements are utilized for the first state update. Next, the updated states are fed into the proposed data association module to establish correspondences between 3D streetlight clusters and 2D detections. The matches will be used for the second state update. More 2D-3D streetlight matches are then obtained through the binary segmentation in the match extension module and are applied to the third state update. Additionally, a tracking recovery module is proposed to resolve localization failures resulting from insufficient matches.
As for the notations and frames, $(\cdot)^{w}$, $(\cdot)^{c}$, $(\cdot)^{o}$ and $(\cdot)^{b}$ denote the world, camera, odometer and body frames. The IMU frame coincides with the body frame. The first IMU frame is regarded as the world frame and the gravity vector is constant in the world frame $\mathbf{g}^{w}=\begin{bmatrix}0&0&-9.81\end{bmatrix}^{\top}$. $\mathbf{R}^{w}_{b}\in SO(3)$ and $\mathbf{p}^{w}_{b}\in\mathbb{R}^{3}$ are rotation and translation from the world frame to body frame. All velocities in this paper are with respect to the world frame. For example, ${}^{w}\mathbf{v}_{b}\in\mathbb{R}^{3}$ is the velocity of body frame with respect to and represented in the world frame. $\tilde{(\cdot)}$ and $\hat{(\cdot)}$ represent the measurement and the estimate, respectively.
@@ III-A Streetlight Map Generation
We employ an image-based object detection method to filter luminous streetlight points from LiDAR maps or measurements for generating the pre-built streetlight map. For each LiDAR scan, points whose projections fall into the detected streetlight boxes are labeled as “streetlight”. The original LiDAR map is generated by the modified FAST-LIO2111https://github.com/hku-mars/FAST_LIO [19] augmented with a fast and accurate loop closure module [20] and only the map points with “streetlight” labels are retained. The streetlight point map is further filtered and clustered by DBSCAN[21] for grouping 3D points into different streetlight instances. Finally the neat 3D streetlight map can be generated as shown in the bottom of Fig. 1.
@@ III-C IMU Propagation Model
where $\tilde{\bm{\omega}}_{t}$ and $\tilde{\bm{a}}_{t}$ represent measurements of angular velocity and specific acceleration in the IMU frame (i.e. the body frame $b$). $\mathbf{n}_{\omega_{t}}\sim\mathcal{N}(\mathbf{0}_{3\times 1},\mathbf{\Sigma}_{\omega})$ and $\mathbf{n}_{a_{t}}\sim\mathcal{N}(\mathbf{0}_{3\times 1},\mathbf{\Sigma}_{a})$ are the noises. The state dynamics model can be derived with the IMU propagation model [23] as
@@
where $\mathbf{Ad}_{\hat{\mathbf{X}}_{t},\hat{\mathbf{b}}_{t}}\triangleq block\_diag(\mathbf{Ad}_{\hat{\mathbf{X}}_{t}},\mathbf{I}_{6\times 6})$ and $\mathbf{Ad}_{\hat{\mathbf{X}}_{t}}$ is the adjoint matrix of $\hat{\mathbf{X}}_{t}$ [16]. $\mathbf{0}$ is the $3\times 3$ zero matrix. The supplementary material[24] gives the detailed derivation. With the above continuous time propagation model and (III-C), the discretized system dynamics can be derived using Euler integration [23]:
@@
where $\mathbf{\Phi}_{t}=\exp_{m}(\mathbf{A}_{t}\Delta t)$ is the state transition matrix and $\exp_{m}(\cdot)$ is the exponential function of matrices. $\hat{\mathbf{P}}_{t}^{+}$ is the updated posterior covariance of $\bm{\xi}_{t}$ at time $t$. $\mathbf{Q}_{t}=block\_diag(\mathbf{\Sigma}_{a\omega},\mathbf{\Sigma}_{a},\mathbf{0},\mathbf{\Sigma}_{b\omega},\mathbf{\Sigma}_{ba})$ denotes the covariance of noise $\mathbf{n}_{t}$.
@@ III-D Odometer Observation Model
The wheel odometer with encoders can provide the linear velocity of the wheel robot. The measured velocity is also assumed to be corrupted by Gaussian white noises
@@
Using the Kalman filtering theory, the Kalman gain $\mathbf{K}_{o_{t}}$, updated states $\hat{\mathbf{X}}_{t}^{+}$ and covariance matrices $\hat{\mathbf{P}}_{t}^{+}$ are obtained:
$\mathbf{K}_{o_{t}}$ $=\hat{\mathbf{P}}_{t}^{-}\mathbf{H}_{o_{t}}^{\top}(\mathbf{H}_{o_{t}}\hat{\mathbf{P}}_{t}^{-}\mathbf{H}_{o_{t}}^{\top}+\hat{\mathbf{R}}_{b_{t}}^{w}\mathbf{\Sigma}_{o}\hat{\mathbf{R}}_{b_{t}}^{w^{\top}})^{-1}\triangleq\begin{bmatrix}\mathbf{K}_{o\xi_{t}}\\ \mathbf{K}_{o\zeta_{t}}\end{bmatrix}$ (24)
$\hat{\mathbf{X}}_{t}^{+}$ $=\exp(\mathbf{K}_{o\xi_{t}}\mathbf{z}_{o_{t}})\hat{\mathbf{X}}_{t}^{-},\quad\hat{\mathbf{b}}_{t}^{+}=\hat{\mathbf{b}}_{t}^{-}+\mathbf{K}_{o\zeta_{t}}\mathbf{z}_{o_{t}}$
$\hat{\mathbf{P}}_{t}^{+}$ $=(\mathbf{I}-\mathbf{K}_{o_{t}}\mathbf{H}_{o_{t}})\hat{\mathbf{P}}_{t}^{-}$ (25)
where $\mathbf{K}_{o_{t}}\in\mathbb{R}^{15\times 3}$ is split into two parts $\mathbf{K}_{o\xi_{t}}\in\mathbb{R}^{9\times 3}$ and $\mathbf{K}_{o\zeta_{t}}\in\mathbb{R}^{6\times 3}$ which are respectively used to update the states on the matrix Lie group and the biases.
@@ III-E1 Data Association
Assuming there are $n_{t}$ detected streetlights and $m_{t}$ 3D streetlight clusters within a certain range of $th_{Lr1}$ from the prior robot pose $\hat{\mathbf{R}}_{b_{t}}^{w},\hat{\mathbf{p}}_{b_{t}}^{w}$. The streetlight observations are represented as $\{B_{t_{i}}=(\tilde{\mathbf{p}}_{t_{i}},\tilde{\bm{l}}_{t_{i}})\}_{i=1}^{n_{t}}$ where $\tilde{\mathbf{p}}_{t_{i}}=\begin{bmatrix}\tilde{u}_{t_{i}}&\tilde{v}_{t_{i}}\end{bmatrix}^{\top}$ is the image coordinate of the center of object detection bounding boxes. $\tilde{\bm{l}}_{t_{i}}=\begin{bmatrix}\tilde{lu}_{t_{i}}&\tilde{lv}_{t_{i}}\end{bmatrix}^{\top}$ contains the width and length of box. The streetlight clusters in the map are represented as $\{L_{j}=(\tilde{\mathbf{C}}_{j}^{w},{{}^{1}}\tilde{\mathbf{P}}^{w}_{j},...,{{}^{J}}\tilde{\mathbf{P}}^{w}_{j}),(-1)\}_{j=1}^{m_{t}+1}$ where $\tilde{\mathbf{C}}_{j}^{w}=\frac{1}{J}\Sigma^{J}_{jj=1}{{}^{jj}}\tilde{\mathbf{P}}^{w}_{j}$ is the geometric center and ${{}^{jj}}\tilde{\mathbf{P}}^{w}_{j}\in\mathbb{R}^{3}$ denotes the 3D coordinate of each point. Since there may be false detections of streetlights, the set of streetlight clusters is extended with element $(-1)$. We score all combinations of matches between 2D and 3D data and the one with the highest score is determined as the correct data association. The score is calculated based on two kinds of match errors.
@@
$s_{t_{proj,ij}}$ $=\mathcal{N}_{3}(\mathbf{r}_{t_{proj,ij}}|\mathbf{0}_{3\times 1},\mathbf{\Sigma_{t_{proj,ij}}})$
@@
where $\mathcal{N}_{3}(\cdot|\mathbf{0}_{3\times 1},\mathbf{\Sigma_{t_{proj,ij}}})$ represents the Gaussian probability density function with mean $\mathbf{0}_{3\times 1}$ and covariance $\mathbf{\Sigma_{t_{proj,ij}}}$ in the 3D Euclidean space. $\mathbf{R}^{c}_{b}$ and $\mathbf{t}_{b}^{c}$ constitute the extrinsics between IMU and camera. $\pi_{c}(\cdot)$ is the perspective projection function of the camera. Angle Error. We also introduce the score based on angle error $\theta_{ij}$ between the back-projection ray of the center of box $B_{t_{i}}$ and the line connecting the center of streetlight cluster $L_{j}$ and the optical center of the camera, which is defined as
$s_{t_{ang,ij}}$ $=\mathcal{N}_{1}(r_{t_{ang,ij}}|0,\sigma^{2}_{t_{ang,ij}})$
@@
We use $d_{ij}$ to indicate whether the $i$-th streetlight observation matches the $j$-th 3D streetlight cluster. $\mathcal{D}_{t}=\{((B_{t_{i}},L_{j})|d_{ij}=1)\}_{i=1}^{n_{t}}$ denotes the possible match combinations. Note that all streetlight observations have a one-to-one mapping with 3D clusters. We expand the set of clusters to $m_{t}+n_{t}$ with $n_{t}$ elements representing no correspondence of 2D observations so that this one-to-one mapping property can be maintained when observation $B_{i}$ is mapped to $(-1)$. To obtain the best combination $\mathcal{D}_{t}^{\star}$, the optimization problem is formulated as the classical assignment problem
@@
After obtaining the 2D-3D matches, the camera observation model can be derived. Only matches of $\mathcal{D}_{t}^{\star}$ without $L_{j}=(-1)$ are selected. For the $i$-th match $(B_{t_{i}},L_{j})$, the observation function is formulated as
@@
where $\mathbf{\Pi}=\begin{bmatrix}\mathbf{I}&\cdots&\mathbf{I}\end{bmatrix}^{\top}\in\mathbb{R}^{3n_{t}\times 3}$. Similar to (24), the Kalman gain, updated states, and covariance can be obtained.
@@ III-E2 Match Extension
III-E3 Degenerated Scenarios
When the streetlights are in a straight line, which is common in the real world, the pose estimation could suffer from degeneration, as shown in Fig. 4. In the degenerated scenarios, the updated pose drifts with time due to the inaccurate prior pose. After the robot experiences a period of degeneration, the matches of new streetlights that are not in a line with already observed streetlights are difficult to search as shown in Fig. 4. We further find the positional drift in $z$-axis and rotational drifts in roll and pitch are more obvious (Section IV-C). The reason is that the odometer can provide 2D motion measurements thus constraining the 3-DoF ($x$, $y$, and yaw). This case results in the projected points of the newly observed streetlights typically being located directly above or below the detection box. Taking advantage of this property, a direct and effective way is proposed. When the degeneration ends (i.e., new streetlights not in a line with already observed ones are detected), each projected center of newly observed streetlight clusters is assigned a rectangle. If the rectangle has an intersection area with detected boxes, the box closest to the projected center is chosen as the match of the corresponding streetlight cluster. Then the pose can be corrected by adding streetlights not in a straight line. Note that the height of the rectangle is significantly larger than its width due to the significant drifts in $z$-axis, roll, and pitch.
Figure 4: Degeneration caused by streetlights in a straight line. The left image shows that all positions on the blue circle satisfy the identical camera observation constraints. An inaccurate prior pose leads to deviation from the true pose. The right images show that the streetlight matches are difficult to search after the degeneration cases.
@@ III-F Tracking Recovery
where the updated pose $(\hat{\mathbf{R}}_{b_{re}}^{w+},\hat{\mathbf{p}}^{w+}_{b_{re}})$ is calculated based on the combination. $neg(i)=1$ when the $i$-th box has no correspondence, otherwise $neg(i)=0$. $\gamma_{1},\gamma_{2},\gamma_{3}$ are utilized for balancing different items. We only use the yaw angle of relative rotation ($yaw(\cdot)$) and the $x$, $y$ components of relative translation ($(\cdot)_{1,2}$). Once there is a combination whose score is less than $th_{RE\_sc}$ and more than $2$ matches are found, it is considered to successfully relocalize the robot. Subsequently, the updated pose of the combination with the minimum score is regarded as the recovered pose.

### 03 Mar 2024 v2

@@ Abstract
Vision-aided localization for low-cost mobile robots in diverse environments has attracted widespread attention recently. Although many current systems are applicable in daytime environments, nocturnal visual localization is still an open problem owing to the lack of stable visual information. An insight from most nocturnal scenes is that the static and bright streetlights are reliable visual information for localization. Hence we propose a nocturnal vision-aided localization system in streetlight maps with a novel data association and matching scheme using object detection methods. We leverage the Invariant Extended Kalman Filter (InEKF) to fuse IMU, odometer, and camera measurements for consistent state estimation at night. Furthermore, a tracking recovery module is also designed for tracking failures. Experimental results indicate that our proposed system achieves accurate and robust localization with less than $0.2\%$ relative error of trajectory length in four nocturnal environments.
@@ I Introduction
  • To mitigate potential errors in nocturnal scenes, the Invariant Extended Kalman Filter (InEKF) is introduced for consistent and stable state estimation.
@@ II System Overview
Fig. 2 shows the framework of the proposed nocturnal localization system. The streetlight map is constructed using only the 3D streetlight points. We introduce the right InEKF due to its accuracy and consistency in the system for fusing monocular camera, IMU, and odometer measurements. Specifically, the robot state is propagated using IMU measurements, and the odometer measurements are utilized for the first state update. Next, the updated states are fed into the proposed data association module to establish correspondences between 3D streetlight clusters and 2D detections. The matches will be used for the second state update. More 2D-3D streetlight matches are then obtained through the binary segmentation in the match extension module and are applied to the third state update. Additionally, a tracking recovery module is proposed to resolve localization failures resulting from insufficient matches.
As for the notations and frames, $(\cdot)^{w}$, $(\cdot)^{c}$, $(\cdot)^{o}$ and $(\cdot)^{b}$ denote the world, camera, odometer and body frames. The IMU frame coincides with the body frame. The first IMU frame is regarded as the world frame and the gravity vector $\mathbf{g}^{w}$ is constant in the world frame. $\mathbf{R}^{w}_{b}\in SO(3)$ and $\mathbf{p}^{w}_{b}\in\mathbb{R}^{3}$ are rotation and translation from the world frame to body frame. All velocities in this paper are with respect to the world frame. For example, ${}^{w}\mathbf{v}_{b}\in\mathbb{R}^{3}$ is the velocity of body frame with respect to and represented in the world frame. $\tilde{(\cdot)}$ and $\hat{(\cdot)}$ represent the measurement and the estimate, respectively.
@@ III-A Streetlight Map Generation
We employ an image-based object detection method[15] to filter luminous streetlight points from LiDAR maps or measurements for generating the pre-built streetlight map. For each LiDAR scan, points whose projections fall into the detected streetlight boxes are labeled as “streetlight”. The original LiDAR map is generated by the modified FAST-LIO2111https://github.com/hku-mars/FAST_LIO [16] augmented with a fast and accurate loop closure module [17] and only the map points with “streetlight” labels are retained. The streetlight point map is further filtered and clustered by DBSCAN[18] for grouping 3D points into different streetlight instances. Finally the neat 3D streetlight map can be generated as shown in the bottom of Fig. 1.
@@ III-C IMU Propagation Model
where $\tilde{\bm{\omega}}_{t}$ and $\tilde{\bm{a}}_{t}$ represent measurements of angular velocity and specific force in the IMU frame (i.e. the body frame $b$). $\mathbf{n}_{\omega_{t}}\sim\mathcal{N}(\mathbf{0}_{3\times 1},\mathbf{\Sigma}_{\omega})$ and $\mathbf{n}_{a_{t}}\sim\mathcal{N}(\mathbf{0}_{3\times 1},\mathbf{\Sigma}_{a})$ are the noises. The state dynamics model can be derived with the IMU propagation model [22] in continuous-time as
@@
where $\mathbf{Ad}_{\hat{\mathbf{X}}_{t},\hat{\mathbf{b}}_{t}}\triangleq block\_diag(\mathbf{Ad}_{\hat{\mathbf{X}}_{t}},\mathbf{I}_{6\times 6})$ and $\mathbf{Ad}_{\hat{\mathbf{X}}_{t}}$ is the adjoint matrix of $\hat{\mathbf{X}}_{t}$ [19]. $\mathbf{0}$ is the $3\times 3$ zero matrix. The supplementary material[23] gives the detailed derivation. With the above continuous-time propagation model and (III-C), the discretized system dynamics can be derived using Euler integration [22]:
@@
where $\mathbf{\Phi}_{t}=\exp_{m}(\mathbf{A}_{t}\Delta t)$ is the state transition matrix, $\exp_{m}(\cdot)$ is the exponential map of matrices. $\hat{\mathbf{P}}_{t}^{+}$ is the posterior covariance of $\bm{\xi}_{t}$ at time $t$. $\mathbf{Q}_{t}=block\_diag(\mathbf{\Sigma}_{\omega},\mathbf{\Sigma}_{a},\mathbf{0},\mathbf{\Sigma}_{b\omega},\mathbf{\Sigma}_{ba})$ is the covariance of $\mathbf{n}_{t}$.
@@ III-D Odometer Observation Model
The wheel odometer with encoders can provide linear velocity measurements for wheel robots. The measured velocity is also assumed to be corrupted by Gaussian white noises
@@
Using the Kalman filtering theory[20], the updated states and covariance can be obtained.
@@ III-E1 Data Association
Assuming there are $n_{t}$ detected streetlights and $m_{t}$ 3D streetlight clusters within a certain range of $th_{Lr1}$ from the robot pose estimate $\hat{\mathbf{R}}_{b_{t}}^{w},\hat{\mathbf{p}}_{b_{t}}^{w}$. The streetlight observations are represented as $\{B_{t_{i}}=(\tilde{\mathbf{p}}_{t_{i}},\tilde{\bm{l}}_{t_{i}})\}_{i=1}^{n_{t}}$ where $\tilde{\mathbf{p}}_{t_{i}}=\begin{bmatrix}\tilde{u}_{t_{i}}&\tilde{v}_{t_{i}}\end{bmatrix}^{\top}$ is the image coordinate of the center of object detection bounding boxes. $\tilde{\bm{l}}_{t_{i}}=\begin{bmatrix}\tilde{lu}_{t_{i}}&\tilde{lv}_{t_{i}}\end{bmatrix}^{\top}$ contains the width and length of box. The streetlight clusters in the map are represented as $\{L_{j}=(\tilde{\mathbf{C}}_{j}^{w},{{}^{1}}\tilde{\mathbf{P}}^{w}_{j},...,{{}^{J}}\tilde{\mathbf{P}}^{w}_{j}),(-1)\}_{j=1}^{m_{t}+1}$ where $\tilde{\mathbf{C}}_{j}^{w}=\frac{1}{J}\Sigma^{J}_{jj=1}{{}^{jj}}\tilde{\mathbf{P}}^{w}_{j}$ is the geometric center and ${{}^{jj}}\tilde{\mathbf{P}}^{w}_{j}\in\mathbb{R}^{3}$ denotes the 3D coordinate of each point. Since there may be false detections of streetlights, the set of streetlight clusters is extended with element $(-1)$. We score all combinations of matches between 2D and 3D data and the one with the highest score is determined as the correct data association. The score is calculated based on two kinds of match errors.
@@
$s_{t_{proj,ij}}$ $=\mathcal{N}(\mathbf{r}_{t_{proj,ij}}|\mathbf{0}_{3\times 1},\mathbf{\Sigma_{t_{proj,ij}}})$
@@
where $\mathcal{N}(\cdot|\mathbf{0}_{3\times 1},\mathbf{\Sigma_{t_{proj,ij}}})$ represents the Gaussian probability density function with mean $\mathbf{0}_{3\times 1}$ and covariance $\mathbf{\Sigma_{t_{proj,ij}}}$ in the 3D Euclidean space. $\mathbf{R}^{c}_{b}$ and $\mathbf{t}_{b}^{c}$ constitute the extrinsics between IMU and camera. $\pi_{c}(\cdot)$ is the perspective projection function of the camera. Angle Error. We also introduce the score based on angle error $\theta_{ij}$ between the back-projection ray of the center of box $B_{t_{i}}$ and the line connecting the center of streetlight cluster $L_{j}$ and the optical center of the camera, which is defined as
$s_{t_{ang,ij}}$ $=\mathcal{N}(r_{t_{ang,ij}}|0,\sigma^{2}_{t_{ang,ij}})$
@@
We use $d_{ij}$ to indicate whether the $i$-th streetlight observation matches the $j$-th 3D streetlight cluster. $\mathcal{D}_{t}=\{((B_{t_{i}},L_{j})|d_{ij}=1)\}_{i=1}^{n_{t}}$ denotes the possible match combinations. Note that all streetlight observations have a one-to-one mapping with 3D clusters. We expand the set of clusters to $m_{t}+n_{t}$ with $n_{t}$ elements representing no correspondence of 2D observations so that this one-to-one mapping feature can be maintained when observation $B_{i}$ is mapped to $(-1)$. To obtain the best combination $\mathcal{D}_{t}^{\star}$, the optimization problem is formulated as the classical assignment problem
@@
After obtaining the 2D-3D matches, the camera observation model can be derived. Only matches of $\mathcal{D}_{t}^{\star}$ without $L_{j}=(-1)$ are selected. For the $i$-th match $(B_{t_{i}},L_{j})$, the observation model is formulated as
@@
where $\mathbf{\Pi}=\begin{bmatrix}\mathbf{I}&\cdots&\mathbf{I}\end{bmatrix}^{\top}\in\mathbb{R}^{3n_{t}\times 3}$.
@@ III-E2 Match Extension
III-E3 Environmental Degeneration
When the streetlights are in a straight line, as shown in Fig. 4, any camera pose lying on the blue dotted circle satisfies the geometric constraints, thus the pose can drift with time. After the robot experiences a period of degeneration, the matches of new streetlights that are not in a line with already observed streetlights are difficult to search as shown in Fig. 4. We further find the positional drift in $z$-axis and rotational drifts in roll and pitch are more obvious (Section IV-C). The reason is that the odometer can provide 2D motion measurements thus constraining the 3-DoF ($x$, $y$, and yaw). This case results in the projected points of the newly observed streetlights typically being located directly above or below the detection box. Using this feature, a direct and effective way is proposed. When the degeneration ends (i.e., new streetlights not in a line with already observed ones are detected), each projected center of newly observed streetlight clusters is assigned a rectangle. If the rectangle has an intersection area with detected boxes, the box closest to the projected center is chosen as the match of the corresponding streetlight cluster. Then the pose can be corrected by adding streetlights not in a straight line. Note that the height of the rectangle is significantly larger than its width due to the significant drifts in $z$-axis, roll, and pitch.
Figure 4: Degeneration case. The left image shows that all positions on the blue circle satisfy the identical camera observation constraints. An inaccurate prior pose leads to deviation from the true pose. The right images show that the streetlight matches are difficult to search after the degeneration cases.
@@ III-F Tracking Recovery
where the updated pose $(\hat{\mathbf{R}}_{b_{re}}^{w+},\hat{\mathbf{p}}^{w+}_{b_{re}})$ is calculated based on the combination. $neg(i)=1$ when the $i$-th box has no correspondence, otherwise $neg(i)=0$. $\gamma_{1},\gamma_{2},\gamma_{3}$ are utilized for balancing different terms. We only use the yaw angle of relative rotation ($yaw(\cdot)$) and the $x$, $y$ components of relative translation ($(\cdot)_{1,2}$). Once there is a combination whose score is less than $th_{RE\_sc}$ and more than $2$ matches are found, it is considered to successfully relocalize the robot. Subsequently, the updated pose of the combination with the minimum score is regarded as the recovered pose.